# De Finetti’s problem with fixed transaction costs and regime switching

### v1

@@ 1. Introduction
As a prototype for SDEs with discontinuous coefficients, it is also interesting to investigate various properties of the solution to (2.1). The transition density of solution $X$ to (2.1) with $\sigma_{+}=\sigma_{-}$ is found in [20] and applied to compute the optimal expected costs in the classical control problem treated in [8]. The above transition density is also used in [11] to identify limiting distribution arising from the central limit theorem, under nonlinear expectation, for random variables with the same conditional variance but ambiguous means. Similarly, the transition density of solution $X$ to (2.1) with $\mu_{+}=0=\mu_{-}$, called oscillating Brownian motion, is found in [21], and used in [12] for computations concerning the central limit theorem, again under nonlinear expectation, for random variables with mean 00 and varying conditional variance. More recently, an explicit expression for transition density of process $X$ in (2.1) is obtained in [13] using solution to the exit problem together with a perturbation approach, which can be applied to express the value function for the control problem in [23].
We focus on a Brownian motion-driven surplus process, as defined in (2.1), where the drift and diffusion coefficients vary depending on whether the process is above or below a fixed threshold. In most of the literature, spatially homogeneous processes, such as Brownian motion and Lévy processes, are used, often resulting in the optimality of a simple barrier-type strategy. In contrast, models involving processes with dynamics dependent on their current values are limited and typically yield non-analytical results. Despite the straightforward dynamics of our surplus process, it has practical applications: it offers a way to model ”regime-switching,” where the regime changes depending on whether the surplus process is above or below the threshold. This approach is particularly suitable for modeling non-stationary premium rates and volatility that can vary based on the company’s financial status. Our regime-switching model differs from the classical regime-switching models in the literature, as it is caused endogenously, while the latter is driven by exogenous factors due to transitions or adjustments in the economic system; see [2], [3], [25] and [27]. Works studying exogenous regime-switching involved optimal dividend problems can be found in [4], [19] and [29]. A work that studies an endogenous regime-switching involved optimal dividend problem appears in [28].
@@
It is important to note that, to the best of the authors’ knowledge, all existing contributions on de Finetti’s optimal dividend problem utilize spatially homogeneous processes or diffusion processes with regular drift and volatility coefficients as their uncontrolled reserve processes. The combination of non-regular drift and volatility coefficients with fixed transaction costs makes the problem particularly complex and demanding. To illustrate this, the optimal selections of $z_{1}$ and $z_{2}$ are interdependent and, in this case, also influenced by the drift/diffusion change-trigger barrier $a$. There are scenarios where both $z_{1}$ and $z_{2}$ are either above or below $a$, as well as cases where $z_{2}$ is above $a$ while $z_{1}$ is below $a$. Each scenario requires a different analysis and approach. To tackle this problem, we first solve the exit problem for (2.1) using a martingale approach, which allows us to derive explicit expressions for the expected dividend function under each $(z_{1},z_{2})$-strategy. Subsequently, we establish sufficient conditions for optimality. This is followed by a case-by-case analysis, demonstrating that by appropriately selecting the barriers, the candidate value function solves the quasi-variational inequality (QVI). As a result, we obtain explicit optimal strategies for all different parameter choices of $\mu_{\pm}$ and $\sigma_{\pm}$ in the model, which aids in better understanding and analyzing the connections between the optimal strategies and these parameters.
@@
The rest of the paper is organized as follows. Section 2 formulates the problem, provides preliminary results, and introduces the two-barrier $(z_{1},z_{2})$-impulsive dividend strategy. In Section 3, we present a complete and explicit characterization of the optimal strategy among the class of $(z_{1},z_{2})$-impulsive dividend strategy. Section 4 is devoted to characterizing the optimal strategy to the targeted dividend control problem. Some lengthy and technical proofs are provided in Appendix A.
@@ 2. Problem formulation and preliminary results
$\operatorname{d}\!X_{t}=(\sigma_{+}\mathbf{1}_{\{X_{t}>a\}}+\sigma_{-}\mathbf{1}_{\{X_{t}\leq a\}})\operatorname{d}\!B_{t}+(\mu_{+}\mathbf{1}_{\{X_{t}>a\}}+\mu_{-}\mathbf{1}_{\{X_{t}\leq a\}})\operatorname{d}\!t,\quad t\geq 0,$ (2.1)
with two-valued drift and two-valued diffusion coefficients. The existence and uniqueness of a strong solution to SDE (2.1) is guaranteed by Theorem 1.3 in Page 55 of [22]. We use it to describe the surplus process of a company before paying dividend. Since we are interested in the process until it ruins, we assume $a>0$, for otherwise, the process before its ruin time follows the classical drifted Brownian motion model. Clearly, one can deal with the case $a\leq 0$ by shifting the process $X$ properly, so our model covers all values of $a\in\mathbb{R}$. Of course, the corresponding ruin time shall be redefined.
@@
$0\leq\sup_{\pi\in\Pi}V_{\pi}(x)\leq x+\frac{\sqrt{\mu_{+}^{2}+2q\sigma_{+}^{2}}+\mu_{+}}{2q\sigma_{+}^{2}}+\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}}{2q\sigma_{-}^{2}},\quad x\geq 0$
@@ 2.1. Verification lemma
$\Pi_{0}:=\big{\{}\pi=(L^{\pi}_{t})_{t\geq 0}\in\Pi;\text{ for any}t\geq 0,\triangle L_{t}^{\pi}\geq\beta\text{ if and only if }\triangle L_{t}^{\pi}>0\big{\}},$ (2.5)
@@
$\mathcal{A}f(x)=\frac{1}{2}(\sigma^{2}_{+}\mathbf{1}_{\{x>a\}}+\sigma^{2}_{-}\mathbf{1}_{\{x\leq a\}})f^{\prime\prime}(x)+(\mu_{+}\mathbf{1}_{\{x>a\}}+\mu_{-}\mathbf{1}_{\{x\leq a\}})f^{\prime}(x),$
@@
By this result, our problem now reduces to find an impulsive strategy $\hat{\pi}\in\Pi_{0}$ (see (2.5)) that fulfills the requirement of Lemma 2.2. We conjecture that the optimal strategy solving the optimal control problem (2.4) shall be some two-barrier impulsive strategy. To show this, we introduce this kind of dividend payout strategy in the subsequent section.
@@ 2.2. Two-barrier strategies and preliminary results
$\begin{cases}\quad L_{t}^{z_{1},z_{2}}=\sum_{0\leq s\leq t}(U_{s-}^{z_{1},z_{2}}-z_{1})\mathbf{1}_{\{U_{s-}^{z_{1},z_{2}}\geq z_{2}\}},\quad t\geq 0,\\ \;\operatorname{d}\!U^{z_{1},z_{2}}_{t}=(\mu_{+}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}>a\}}+\mu_{-}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}\leq a\}})\operatorname{d}\!t-(U_{t-}^{z_{1},z_{2}}-z_{1})\mathbf{1}_{\{U_{t-}^{z_{1},z_{2}}\geq z_{2}\}}\\ \qquad\qquad\quad+(\sigma_{+}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}>a\}}+\sigma_{-}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}\leq a\}})\operatorname{d}\!B_{t},\quad t\geq 0,\\ \quad\!U^{z_{1},z_{2}}_{0-}=x.\end{cases}$
@@
In the following, we aim to find an explicit expression for $V^{z_{2}}_{z_{1}}(x)$ so that we can apply Lemma 2.2 to derive an optimal strategy to (2.4). To this end, we first define two functions $\𝑎g^{\pm}\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$ that satisfy the following ordinary differential equation (ODE) except at $a$:
@@
Then, the following two functions $\𝑎g^{\pm}\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$ satisfy ODE (2.6) (except at $a$):
@@
Then, one has $\𝑎g\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$ and it satisfies ODE (2.6) (except at $a$).
@@
$\mathbb{E}_{x}[e^{-qT_{y}}\mathbf{1}_{\{T_{y}<T_{z}\}}]=\frac{g^{+}(z)g^{-}(x)-g^{-}(z)g^{+}(x)}{g^{-}(y)g^{+}(z)-g^{-}(z)g^{+}(y)},$
@@
$\mathbb{E}_{x}[e^{-qT_{z}}\mathbf{1}_{\{T_{z}<T_{y}\}}]=\frac{g^{+}(y)g^{-}(x)-g^{-}(y)g^{+}(x)}{g^{-}(z)g^{+}(y)-g^{-}(y)g^{+}(z)}.$
@@
Vz1z2(x)={g(z1)(z2z1β)g(z2)g(z1)+xz1β,xz2,
g(x)(z2z1β)g(z2)g(z1),0x<z2.
[…]
@@
$\mathcal{D}_{\zeta}$ :=$:=\{(x,y)\in[0,\infty)^{2}:x+\beta\leq y\},$
@@
Indeed, we will prove that the set $\mathcal{M}_{\zeta}$ consists of either one, two or three elements in Theorems 3.1-3.4.
@@
Vz1z2(x)={g(z2)g(z2)+xz2,xz2,
g(x)g(z2),0x<z2.
[…]
@@
$\mathcal{N}^{+}$ :=$:=\{(z_{1},z_{2}):0<z_{1}<z_{2}<\infty,\,\psi(z_{1},z_{2})=\beta,g^{\prime}(z_{1})=g^{\prime}(z_{2})\},$
$\mathcal{N}^{-}$ :=$:=\{(0,z_{2}):0<z_{2}<\infty,\,\psi(0,z_{2})=\beta\}.$
@@ 3. Explicit characterization of $\mathcal{M}_{\zeta}$
This section is devoted to the characterization of the explicit form of the set $\mathcal{M}_{\zeta}$ corresponding to four mutually exclusive and collectively exhaustive cases: (1) $\mu_{\pm}>0$; (2) $\mu_{\pm}\leq 0$; (3) $\mu_{+}\leq 0$ and $\mu_{-}>0$; and, (4) $\mu_{+}>0$ and $\mu_{-}\leq 0$. We also provide, in Subsection 3.5, several general properties of $\mathcal{M}_{\zeta}$ that can help to better understand and analyze the connections between the optimal dividend strategy and the model parameters.
@@
whenever they are well-defined (note that $\ln x$ is not well-defined for $x\leq 0$).
@@
Recall that, under Case (i), $g^{\prime}(x)$ is continuous, strictly decreasing and continuously differentiable on $(0,a)$ and strictly increasing and continuously differentiable on $(a,\infty)$ by Proposition 3.1. Let $(g^{\prime})_{-}^{-1}(x):=\inf\{z\in[0,a];g^{\prime}(z)\leq x\}$ for $x\in[g^{\prime}(a),\infty)$ and $(g^{\prime})_{+}^{-1}(x)$ be the inverse function of $[a,\infty)\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(a),\infty)$. Define $a_{4}:=(g^{\prime})^{-1}_{+}(g^{\prime}(0))$. Further, define the unary function $\phi$ as
@@
Under Case (iv), $g^{\prime}(x)$ is continuous, strictly decreasing (resp., increasing) and continuously differentiable on $(0,a_{1})$ and $(a,x_{0})$ (resp., $(a_{1},a)$ and $(x_{0},\infty)$). Let $(g^{\prime})_{1}^{-1}$ be defined the same as $(g^{\prime})_{-}^{-1}$ but with $a_{1}$ in place of $a$; $(g^{\prime})_{2}^{-1}$ be the inverse function of $[a_{1},a]\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(a_{1}),g^{\prime}(a)]$; $(g^{\prime})_{3}^{-1}$ be the inverse function of $[a,x_{0}]\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(x_{0}),g^{\prime}(a)]$; and, $(g^{\prime})_{4}^{-1}$ be the inverse function of $[x_{0},\infty)\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(x_{0}),\infty)$. Denote $a_{5}:=\inf\{x\geq x_{0};g^{\prime}(x)\geq g^{\prime}(a_{1})\}$ and $a_{6}:=(g^{\prime})^{-1}_{4}(g^{\prime}(a))$. Furthermore, put
@@
$\omega_{2}(x)$ :=$:=\begin{cases}\int_{(g^{\prime})_{3}^{-1}(g^{\prime}(x))}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{0},x_{1}),\\ \int_{(g^{\prime})_{1}^{-1}(g^{\prime}(x))}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{1},\infty).\end{cases}$ (3.7)
@@
    $\mathcal{M}_{\zeta}=\begin{cases}\{(\tilde{z}_{1},\tilde{z}_{2})\},&\text{if $\beta\in A_{1}\cup A_{3}$, }\\ \{(\bar{z}_{1},\bar{z}_{2})\},&\text{if $\beta\in A_{2}\cap\overline{A}_{3}$,}\\ \{(\tilde{z}_{1},\tilde{z}_{2})\}\cup\{(\bar{z}_{1},\bar{z}_{2})\},&\text{otherwise,}\end{cases}$
    where $(\tilde{z}_{1},\tilde{z}_{2}):=((g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),\omega_{1}^{-1}(\beta))$, $(\bar{z}_{1},\bar{z}_{2}):=((g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),\omega_{2}^{-1}(\beta))$, $A_{1}:=\{\beta>0:g^{\prime}(\omega_{1}^{-1}(\beta))<g^{\prime}(\omega_{2}^{-1}(\beta))\}$, $A_{2}:=\{\beta>0:g^{\prime}(\omega_{1}^{-1}(\beta))>g^{\prime}(\omega_{2}^{-1}(\beta))\}$, $A_{3}:=(\omega_{2}(x_{1}),\infty)$, and $\overline{A}:=(0,\infty)\setminus A$. Note that, when $\beta=\omega_{1}(x_{2})$, one has $\omega_{1}^{-1}(\beta)=\{(g^{\prime})_{2}^{-1}(g^{\prime}(x_{2})),x_{2}\}$, in which case, $\{(\tilde{z}_{1},\tilde{z}_{2})\}$ is understood as $\{(\tilde{z}_{1},(g^{\prime})_{2}^{-1}(g^{\prime}(x_{2}))),(\tilde{z}_{1},x_{2})\}$.
@@
The following Theorem 3.2 explicitly characterizes the set $\mathcal{M}_{\zeta}$ defined by (2.13) as a singleton set. The proof of Theorem 3.2 is similar to that of Theorem 3.1 and hence omitted.
@@
Suppose that $\mu_{\pm}\leq 0$. The set $\mathcal{M}_{\zeta}$ is singleton and given as

### v2

@@ 1. Introduction
As a prototype for SDEs with discontinuous coefficients, it is also interesting to investigate various properties of the solution to (2.1). The transition density of solution $X$ to (2.1) with $\sigma_{+}=\sigma_{-}$ is found in [20] and applied to compute the optimal expected costs in the classical control problem treated in [8]. The above transition density is also used in [11] to identify limiting distribution arising from the central limit theorem, under nonlinear expectation, for random variables with the same conditional variance but ambiguous means. Similarly, the transition density of solution $X$ to (2.1) with $\mu_{+}=0=\mu_{-}$, called oscillating Brownian motion, is found in [21], and used in [12] for computations concerning the central limit theorem, again under nonlinear expectation, for random variables with mean 0 and varying conditional variance. More recently, an explicit expression for transition density of process $X$ in (2.1) is obtained in [13] using solution to the exit problem together with a perturbation approach, which can be applied to express the value function for the control problem in [23].
We focus on a Brownian motion-driven surplus process, as defined in (2.1), where the drift and diffusion coefficients vary depending on whether the process is above or below a fixed threshold. In most of the literature, spatially homogeneous processes, such as Brownian motion and Lévy processes, are used, often resulting in the optimality of a simple barrier-type strategy. In contrast, models involving processes with dynamics dependent on their current values are limited and typically yield non-analytical results. Despite the straightforward dynamics of our surplus process, it has practical applications: it offers a way to model “regime-switching”, where the regime changes depending on whether the surplus process is above or below the threshold. This approach is particularly suitable for modeling non-stationary premium rates and volatility that can vary based on the company’s financial status. Our regime-switching model differs from the classical regime-switching models in the literature, as it is caused endogenously, while the latter is driven by exogenous factors due to transitions or adjustments in the economic system; see [2], [3], [25] and [27]. Works studying exogenous regime-switching involved optimal dividend problems can be found in [4], [19] and [29]. A work that studies an endogenous regime-switching involved optimal dividend problem appears in [28].
@@
It is important to note that, to the best of the authors’ knowledge, all existing contributions on de Finetti’s optimal dividend problem utilize spatially homogeneous processes or diffusion processes with regular drift and volatility coefficients as their uncontrolled reserve processes. The combination of non-regular drift and volatility coefficients with fixed transaction costs makes the problem particularly complex and demanding. In such a setting, the optimal selections of $z_{1}$ and $z_{2}$ are interdependent and, in this case, also influenced by the drift/diffusion change-trigger barrier $a$. There are scenarios where both $z_{1}$ and $z_{2}$ are either above or below $a$, as well as cases where $z_{2}$ is above $a$ while $z_{1}$ is below $a$. Each scenario requires a different analysis and approach. To tackle this problem, we first solve the exit problem for (2.1) using a martingale approach, which allows us to derive explicit expressions for the expected dividend function under each $(z_{1},z_{2})$-strategy. Subsequently, we establish sufficient conditions for optimality. This is followed by a case-by-case analysis, demonstrating that by appropriately selecting the barriers, the candidate value function solves the quasi-variational inequality (QVI). As a result, we obtain explicit optimal strategies for all different parameter choices of $\mu_{\pm}$ and $\sigma_{\pm}$ in the model, which aids in better understanding and analyzing the connections between the optimal strategies and these parameters.
The contributions of this paper can be summarized from two perspectives: one pertains to the problem formulation, and the other to the solution methodology. We subsequently provide a detailed elaboration of these contributions.
From a problem formulation perspective, our main contribution is to solve, for the first time, an optimal dividend problem by simultaneously incorporating two critical features: (1) an endogenous state-triggered switching mechanism, and (2) fixed transaction costs, thereby framing the problem as a challenging impulse control problem. While each feature has been studied in isolation, their combination has not been previously addressed. On the one hand, most existing studies model regime switching exogenously using a Markov chain (e.g., [4], [19], [25], [29]). A recent work by [28] considers endogenous regime switching. However, none of these works incorporate fixed transaction costs. On the other hand, many works include fixed costs (e.g., [5], [7], [10]), but their models are spatially homogeneous and exclude regime switching. Others ([6], [24], [17]) allow state-dependent coefficients but impose strong regularity assumptions (e.g., Lipschitz continuity, differentiability). In contrast, our model allows for discontinuous drift and volatility coefficients, which change abruptly at a threshold and take distinct values on either side.
From a methodological perspective, existing methods for dividend problems with regime switching (e.g., [19], [28], [29]) rely on regular or singular control and are not applicable to our impulse control setting with fixed transaction costs. Similarly, methods for fixed-cost problems (e.g., [5], [6], [7], [10], [17], [24]) generally assume either spatial homogeneity, constant coefficients, or regular (smooth) state-dependence coefficients, none of which hold in our framework of discontinuous coefficients. To address these challenges, we develop a novel analytical methodology for solving the associated QVI, which includes the following four technical contributions:
First, through extensive technical analysis, we provide a complete and unified characterization of the piecewise convexity/concavity of the function $g$ (Propositions 3.13.4), identifying 11 mutually exclusive and collectively exhaustive cases (with 21 sub-cases) that fully describe the geometry of $g$ and facilitate the explicit construction of candidate optimal strategies.
Second, we introduce a novel construction-based method to characterize the set of candidate optimal impulse dividend strategies $\mathcal{M}_{\zeta}$ (Theorems 3.13.4), yielding closed-form expressions for all relevant quantities and enabling rigorous sensitivity analysis and numerical implementation.
Third, in the proof of Theorem 4.1, we overcome the challenges posed by discontinuous coefficients by constructing new inequalities (see (4.4) and (4.6), etc.) and proposing three calibrated sufficient (of which the union of the first two is also necessary) conditions to verify optimality.
Finally, to establish the monotonicity result in Proposition 3.6, we employ a novel re-parameterization technique, transforming the analysis from the “cost (i.e., $\beta$)” space to a more tractable “slope of the value function” space. This idea may offer useful insights for other control problems.
@@
The rest of the paper is organized as follows. Section 2 formulates the problem, provides preliminary results, and introduces the two-barrier $(z_{1},z_{2})$-impulsive dividend strategy. In Section 3, we present a complete and explicit characterization of the optimal strategy among the class of $(z_{1},z_{2})$-impulsive dividend strategies. Section 4 is devoted to characterizing the optimal strategy to the targeted dividend control problem. Some lengthy and technical proofs are provided in Appendix A.
@@ 2. Problem formulation and preliminary results
$\operatorname{d}\!X_{t}=(\sigma_{+}\mathbf{1}_{\{X_{t}>a\}}+\sigma_{-}\mathbf{1}_{\{X_{t}\leq a\}})\operatorname{d}\!B_{t}+(\mu_{+}\mathbf{1}_{\{X_{t}>a\}}+\mu_{-}\mathbf{1}_{\{X_{t}\leq a\}})\operatorname{d}\!t,\quad t\geq 0.$ (2.1)
The existence and uniqueness of a strong solution to SDE (2.1) are guaranteed by Theorem 1.3 in Page 55 of [22]. We use it to describe the surplus process of a company before paying dividend. Since we are interested in the process until it ruins, we assume $a>0$, for otherwise, the process before its ruin time follows the classical drifted Brownian motion model. Clearly, one can deal with the case $a\leq 0$ by shifting the process $X$ properly, so our model covers all values of $a\in\mathbb{R}$. Of course, the corresponding ruin time shall be redefined.
@@
$0\leq\sup_{\pi\in\Pi}V_{\pi}(x)\leq x+\frac{\sqrt{\mu_{+}^{2}+2q\sigma_{+}^{2}}+\mu_{+}}{2q}+\frac{\sqrt{\mu_{-}^{2}+2q\sigma_{-}^{2}}+\mu_{-}}{2q},\quad x\geq 0.$ (2.5)
@@ 2.1. Verification lemma
[…]
@@
$\mathcal{A}f(x):=\frac{1}{2}(\sigma^{2}_{+}\mathbf{1}_{\{x>a\}}+\sigma^{2}_{-}\mathbf{1}_{\{x\leq a\}})f^{\prime\prime}(x)+(\mu_{+}\mathbf{1}_{\{x>a\}}+\mu_{-}\mathbf{1}_{\{x\leq a\}})f^{\prime}(x),$
@@
By this result, our problem now reduces to finding an impulsive strategy $\hat{\pi}\in\Pi_{0}$ (see (2.6)) that fulfills the requirement of Lemma 2.2. We conjecture that the optimal strategy solving the optimal control problem (2.4) shall be some two-barrier impulsive strategy. To show this, we introduce this kind of dividend payout strategy in the subsequent section.
@@ 2.2. Two-barrier strategies and preliminary results
$\begin{cases}\quad L_{t}^{z_{1},z_{2}}=\sum_{0\leq s\leq t}(U_{s-}^{z_{1},z_{2}}-z_{1})\mathbf{1}_{\{U_{s-}^{z_{1},z_{2}}\geq z_{2}\}},\quad t\geq 0,\\ \;\operatorname{d}\!U^{z_{1},z_{2}}_{t}=(\mu_{+}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}>a\}}+\mu_{-}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}\leq a\}})\operatorname{d}\!t-(U_{t-}^{z_{1},z_{2}}-z_{1})\mathbf{1}_{\{U_{t-}^{z_{1},z_{2}}\geq z_{2}\}}\\ \qquad\qquad\quad+(\sigma_{+}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}>a\}}+\sigma_{-}\mathbf{1}_{\{U^{z_{1},z_{2}}_{t}\leq a\}})\operatorname{d}\!B_{t},\quad t\geq 0,\\ \quad\!U^{z_{1},z_{2}}_{0-}=x.\end{cases}$ (2.7)
@@
In the following, we aim to find an explicit expression for $V^{z_{2}}_{z_{1}}(x)$ so that we can apply Lemma 2.2 to derive an optimal strategy to (2.4). To this end, we first define two functions $g^{\pm}\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$ that satisfy the following ordinary differential equation (ODE) except at $a$:
@@
Then, the following two functions $g^{\pm}\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$ satisfy ODE (2.8) (except at a):
@@
Then, one has $g\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$ and it satisfies ODE (2.8) (except at a).
@@
$\mathbb{E}_{x}[e^{-qT_{y}}\mathbf{1}_{\{T_{y}<T_{z}\}}]=\frac{g^{+}(z)g^{-}(x)-g^{-}(z)g^{+}(x)}{g^{-}(y)g^{+}(z)-g^{-}(z)g^{+}(y)},$ (2.14)
@@
$\mathbb{E}_{x}[e^{-qT_{z}}\mathbf{1}_{\{T_{z}<T_{y}\}}]=\frac{g^{+}(y)g^{-}(x)-g^{-}(y)g^{+}(x)}{g^{-}(z)g^{+}(y)-g^{-}(y)g^{+}(z)}.$ (2.15)
@@
$V^{z_{2}}_{z_{1}}(x)=\begin{cases}\frac{g(z_{1})(z_{2}-z_{1}-\beta)}{g(z_{2})-g(z_{1})}+x-z_{1}-\beta,&x\geq z_{2},\vskip 12.0pt plus 4.0pt minus 4.0pt\\ \frac{g(x)(z_{2}-z_{1}-\beta)}{g(z_{2})-g(z_{1})},&0\leq x<z_{2}.\end{cases}$ (2.16)
@@
$\mathcal{D}_{\zeta}$ $:=\{(x,y)\in[0,\infty)^{2}:x+\beta\leq y\},$
@@
$V^{z_{2}}_{z_{1}}(x)=\begin{cases}\frac{g(z_{2})}{g^{\prime}(z_{2})}+x-z_{2},&x\geq z_{2},\vskip 12.0pt plus 4.0pt minus 4.0pt\\ \frac{g(x)}{g^{\prime}(z_{2})},&0\leq x<z_{2}.\end{cases}$ (2.19)
@@
$\mathcal{N}^{+}$ $:=\{(z_{1},z_{2}):0<z_{1}<z_{2}<\infty,\,\psi(z_{1},z_{2})=\beta,g^{\prime}(z_{1})=g^{\prime}(z_{2})\},$
$\mathcal{N}^{-}$ $:=\{(0,z_{2}):0<z_{2}<\infty,\,\psi(0,z_{2})=\beta\}.$ (2.23)
@@ 3. Explicit characterization of $\mathcal{M}_{\zeta}$
This section is devoted to the characterization of the explicit form of the set $\mathcal{M}_{\zeta}$ corresponding to four mutually exclusive and collectively exhaustive cases: (1) $\mu_{\pm}>0$; (2) $\mu_{\pm}\leq 0$; (3) $\mu_{+}\leq 0$ and $\mu_{-}>0$; and (4) $\mu_{+}>0$ and $\mu_{-}\leq 0$. We also provide, in Subsection 3.5, several general properties of $\mathcal{M}_{\zeta}$ that can help to better understand and analyze the connections between the optimal dividend strategy and the model parameters.
@@
whenever they are well-defined (note that $\ln x$ is not well-defined for $x\leq 0$). The constants defined in (3.1) and (3.2) are instrumental in establishing the piecewise convexity (concavity) of $g$ (see, Propositions 3.1-3.4), a property that is fundamental to the analysis in this paper.
@@
Recall that, under Case (i), $g^{\prime}(x)$ is continuous, strictly decreasing and continuously differentiable on $(0,a)$ and strictly increasing and continuously differentiable on $(a,\infty)$ by Proposition 3.1. Let $(g^{\prime})_{-}^{-1}(x):=\inf\{z\in[0,a];g^{\prime}(z)\leq x\}$ for $x\in[g^{\prime}(a),\infty)$ and $(g^{\prime})_{+}^{-1}(x)$ be the inverse function of $a,\infty)\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(a),\infty)$. Define $a_{4}:=(g^{\prime})^{-1}_{+}(g^{\prime}(0))$. Further, define the unary function $\phi$ as
@@
Under Case (iv), $g^{\prime}(x)$ is continuous, strictly decreasing (resp., increasing) and continuously differentiable on $(0,a_{1})$ and $(a,x_{0})$ (resp., $(a_{1},a)$ and $(x_{0},\infty)$). Let $(g^{\prime})_{1}^{-1}$ be defined the same as $(g^{\prime})_{-}^{-1}$ but with $a_{1}$ in place of $a$; $(g^{\prime})_{2}^{-1}$ be the inverse function of $a_{1},a]\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(a_{1}),g^{\prime}(a)]$; $(g^{\prime})_{3}^{-1}$ be the inverse function of $a,x_{0}]\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(x_{0}),g^{\prime}(a)]$; and, $(g^{\prime})_{4}^{-1}$ be the inverse function of $x_{0},\infty)\ni x\mapsto g^{\prime}(x)\in[g^{\prime}(x_{0}),\infty)$. Denote $a_{5}:=\inf\{x\geq x_{0};g^{\prime}(x)\geq g^{\prime}(a_{1})\}$ and $a_{6}:=(g^{\prime})^{-1}_{4}(g^{\prime}(a))$. Furthermore, put
@@
$\omega_{2}(x)$ $:=\begin{cases}\int_{(g^{\prime})_{3}^{-1}(g^{\prime}(x))}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{0},x_{1}),\\ \int_{(g^{\prime})_{1}^{-1}(g^{\prime}(x))}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{1},\infty).\end{cases}$ (3.7)
@@
    $\mathcal{M}_{\zeta}=\begin{cases}\{(\tilde{z}_{1},\tilde{z}_{2})\},&\text{if $\beta\in A_{1}\cup A_{3}$, }\\ \{(\bar{z}_{1},\bar{z}_{2})\},&\text{if $\beta\in A_{2}\cap\overline{A}_{3}$,}\\ \{(\tilde{z}_{1},\tilde{z}_{2})\}\cup\{(\bar{z}_{1},\bar{z}_{2})\},&\text{if $\beta\in\overline{A_{1}\cup A_{2}\cup A_{3}}$,}\end{cases}$
    where $(\tilde{z}_{1},\tilde{z}_{2}):=((g^{\prime})_{1}^{-1}(g^{\prime}(\omega_{1}^{-1}(\beta))),\omega_{1}^{-1}(\beta))$, $(\bar{z}_{1},\bar{z}_{2}):=((g^{\prime})_{3}^{-1}(g^{\prime}(\omega_{2}^{-1}(\beta))),\omega_{2}^{-1}(\beta))$, $A_{1}:=\{\beta>0:g^{\prime}(\omega_{1}^{-1}(\beta))<g^{\prime}(\omega_{2}^{-1}(\beta))\}$, $A_{2}:=\{\beta>0:g^{\prime}(\omega_{1}^{-1}(\beta))>g^{\prime}(\omega_{2}^{-1}(\beta))\}$, $A_{3}:=(\omega_{2}(x_{1}),\infty)$, and $\overline{A}:=(0,\infty)\setminus A$. It holds that $A_{2}\cap A_{3}=\emptyset$. We mention that, when $\beta=\omega_{1}(x_{2})$, one has $\omega_{1}^{-1}(\beta)=\{(g^{\prime})_{2}^{-1}(g^{\prime}(x_{2})),x_{2}\}$; in this case, $\{(\tilde{z}_{1},\tilde{z}_{2})\}$ is understood as $\{(\tilde{z}_{1},(g^{\prime})_{2}^{-1}(g^{\prime}(x_{2}))),(\tilde{z}_{1},x_{2})\}$.
@@
The following Theorem 3.2 explicitly characterizes the set $\mathcal{M}_{\zeta}$ defined by (2.17) as a singleton. The proof of Theorem 3.2 is similar to that of Theorem 3.1 and hence omitted.
@@
Suppose that $\mu_{\pm}\leq 0$. The set $\mathcal{M}_{\zeta}$ is a singleton set as